OpenStudy (anonymous):

Find three consecutive odd integers such that the sum of the first two integers is 91 more than the last integer.

OpenStudy (poopsiedoodle):

I tell ya, I just have no idea. Sorry D:

OpenStudy (anonymous):

1.93 95 97 2.16 18 20 3.35 36 37

OpenStudy (anonymous):

consecutive odd integers (x , x + 2, x + 4) sum of 1st two integers = 91 more then the last integer... x + x + 2 = x + 4 + 91 (combine like terms) 2x + 2 = x + 95 2x - x = 95 - 2 x = 93 The other two numbers are : x + 2 = 93 + 2 = 95 x + 4 = 93 + 4 = 97 check... 2x + 2 = x + 95 2(93) + 2 = 93 + 95 186 + 2 = 93 + 95 188 = 188 (correct) Do you have any questions ? Do you understand everything I did ?